Problem 1

  •   A hockey puck has a momentum of 3.8 kgm/s [E]. If its speed is 24 m/s, what is the mass of the puck?

Problem 2

  •  During competitive world-class events, a four-person bobsled experiences an average net force of magnitude 1390 kg•m/s2 during the first 5.0 s of a run.
  • (a) What will be the magnitude of the impulse provided to the bobsled?
  • (b) If the sled has the maximum mass of 630 kg, what will be the speed of the sled?

Problem 3

  • What impulse occurs when an average force of 56 kg•m/s2 is exerted on a vehicle for 10.5 seconds?
1
Expert's answer
2021-05-05T10:54:20-0400

1)

"p=mv,""m=\dfrac{p}{v}=\dfrac{3.8\ kgm/s}{24\ \dfrac{m}{s}}=0.158\ kg."

2) a)

"J=F_{avg}\Delta t=1390\ N\cdot5.0\ s=6950\ Ns."

b)

"F_{avg}\Delta t=m(v_i-v_f)=mv_f,""v_f=\dfrac{F_{avg}\Delta t}{m}=\dfrac{6950\ Ns}{630\ kg}=11\ \dfrac{m}{s}."

3)

"J=F_{avg}\Delta t=56\ N\cdot10.5\ s=588\ Ns."